A big pipe can fill a tank in 7 minutes less than a small pipe. If they work together, the total time taken is 12 minutes. How long will it take for the big pipe to fill a tank alone?
Answer by deborabr(181)

small pipe = x
big pipe = x-7
x+x-7=12
2x=12+7
2x=19
x=19/2
the small pipe takes 19/2=9,5 min alone
big pipe takes 9,5-7 = 2,5 min